a)      Nominal one year interest rates in South Africa and Australia are 6.75% and 2.9% respectively. The current AUD ZAR spot rate is 10.0314. You think that purchasing power parity holds and the one year real interest rate in South Africa is 1.75% and 1% in Australia. What would your expected spot rate for AUD ZAR be?                                                                                           

b)     One year interest rates in Australia and Papua New Guinea are 2.9% and 6.25% respectively. Today’s spot rate is AUD PGK 2.3027. If we assume that the international Fisher effect holds, what would you expect the AUD PGK spot rate to be in one year?
Answer #1

a)

1. Calculation of Inflation Rate:

South Africa Australia
Nominal Rate 6.75% 2.90%
Real Rate 1.75% 1%
Inflation Rate 5.00% 1.90%

As per Purchasing Power Parity Theorem

FS=1+rHC/1+rFC

Where F=Forward Rate/Expected Spot Rate

S = Spot Rate

rHC = Rate of Interest on Home Country (South Africa)

rFC = Rate of Inflation of Foreign Country (Australia)

F=(1+5%) * 10.0314 / 1+1.90%)

F=(10.0314*1.05)/1.019

F=10.3366

i.e. Expected Spot Rate is AUD = ZAR 10.3366